OpenStudy (anonymous):

Write a standard equation of the parabola with focus (2,1) and vertex (-5,1)

OpenStudy (jdoe0001):

 so, the vertex is at (-5, 1) and the focus is at (2, 1) notice the "y" didn't change, the "x" did that means the parabola is opening towards the focus, that is, to the right-handside as show above

OpenStudy (anonymous):

so what would the equation be for that in standard form

OpenStudy (jdoe0001):

$$ (y-\color{blue}{k})^2 = 4p(x-\color{red}{h})\\ \text{"p" is the distance from the vertex to the focus}\\ \text{from -5 to +2, you have 7 units, thus p = 7}\\ \text{the vertex is at} (\color{red}{h}, \color{blue}{k}) $$

OpenStudy (jdoe0001):

so, you have already (h, k) = (-5, 1) and p = 7

OpenStudy (anonymous):

Thank you so much and do uhappen to know how to do this And do you happen to know how you would write this equation into a calculater y^2-4x+2y+5=0

OpenStudy (jdoe0001):

sadly, most calculators use function based on "y" for their plotting, so for y^2-4x+2y+5=0, you'd need to solve for "y", which can get cumbersome I just found out hehe

OpenStudy (jdoe0001):

but what you can do is, make that into a (y-k)^2 = 4p(x-h) form, gimme a few secs

OpenStudy (jdoe0001):

\(y^2+2y = 4x-5\) # completing the square for the perfect trinomial on the left-hand side \((y^2+2y+1^2) = 4x-5\) # getting the perfect square trinomial, and grouping on the right-side \((y+1)^2 = 4(x-\frac{5}{4})\)

OpenStudy (jdoe0001):

so, your vertex is at (5/4, -1) to plot the parabola opening, it has a y^2 component, meaning is opens horizontally 4P = 4, meaning P = 1, is a positive value, meaning it opens to the right-hand side so just pick any 2 points greater than 5/4 for "x" and solve for "y"

OpenStudy (jdoe0001):

so, when say x = 3 \((y+1)^2 = 7, y = \sqrt{7} -1 = 1.645\)

OpenStudy (jdoe0001):

though once you have the equation above, you can just also give a calculator \((y+1)^2 = 4\pmatrix{x-\frac{5}{4}} \implies y = \sqrt{4\pmatrix{x-\frac{5}{4}}}-1 \)

OpenStudy (jdoe0001):

I just notice one tiny mistake I made, when doing the completing of the square, I was supposed to do -1 to the left side, so the correct equation will be $$\large{ (y+1)^2 = 4\pmatrix{x-\frac{6}{4}} \implies y = \sqrt{4\pmatrix{x-\frac{3}{2}}}-1 } $$